# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v4

@@ 4 Application of Model Merging in Different Machine Learning Subfields
Table 3: A summary of the application of model merging techniques in different machine learning subfields.
@@
4.1 Model Merging in Continual Learning
@@
Continual Learning (CL) involves training a model using a streaming, non-stationary data stream. The primary challenge in CL is the ‘catastrophic forgetting’ [155, 203] problem; that is, the CL model’s prediction accuracy for old tasks drops dramatically after training on new tasks. The mainstream CL methods are mainly divided into memory replay-based methods, architecture expansion-based methods, regularization-based methods, and subspace projection-based methods [203]. Recently, there has been a growing interest in using model merging to address the catastrophic forgetting problem. This novel approach offers several benefits, such as avoiding additional parameters and inference costs associated with network expansion-based methods and eliminating the need to cache old data as required by memory-based methods.
4.1.1 Model Merging to Mitigate Catastrophic Forgetting
@@
To overcome catastrophic forgetting and maintain new learning ability, many works have proposed to merge models for old tasks and update models for new tasks [209, 161, 138]. Tangent Model Composition [117] proposes fine-tuning each task independently in the tangent space of the pre-trained model and then linearly fine-tuning these models to perform CL. This approach does not depend on the specific settings of CL and can be easily applied to task, class, and domain-incremental learning scenarios. In addition, ITA [143] emphasizes the necessity for the fine-tuned model to be in the same basin as the pre-trained model to ensure the composability of nonlinear models. It introduces a regularization term similar to EWC [94] in traditional CL to constrain the distance between the fine-tuned weights and the pre-trained weights when training the independent model. WARP [152] suggests linearly interpolating the pre-trained LLM’s weights with its aligned weights via RLHF on a preference dataset, thus mitigating the forgetting of knowledge from the pre-trained LLM. BAM [7] continuously adapts LLMs to new languages by merging models while preserving general capabilities. Model Tailor [243] explores the problem of catastrophic forgetting during fine-tuning of MLLMs, and proposes to merge only the most important subset of parameters in the fine-tuned MLLM model into the pre-trained MLLM model, so as to retain the generalization ability of the pre-trained model as much as possible, while compensating the selected weights to reduce the performance of the fine-tuning task. MagMax [125] merges pruned task vectors to further alleviate parameter sign conflicts and old knowledge forgetting. Equifinality, PAINT [70] and LM-Cocktail [213] interpolate the weights of the fine-tuned model and the zero-shot model to improve accuracy on downstream tasks without degrading accuracy on supported/general tasks.
@@
In contrast to merging full models, some research focuses on merging parameter-efficient modules. Chitale et al. [22] propose a CL method based on task arithmetic [71]. This method first fine-tunes a task-specific LoRA for each task, then constructs a task vector based on the difference between fine-tuned and pre-trained models. Multiple task vectors are then merged, and a small amount of data (10 samples per class) is used to fine-tune the merged model. Compared to traditional CL methods, particularly those based on replay, this approach eliminates the need to replay data from old tasks at each iteration, thereby accelerating model training. Additionally, fine-tuning the merged model with a class-balanced subset helps mitigate CL model bias. Similarly, DynaMMo [144] applies lightweight model merging (i.e., Adapter) in a CL setting for medical images. In contrast to architecture expansion-based CL methods, this approach does not result in a linear increase in the number of parameters with the number of tasks. Unlike the static aggregated parameter-efficient fine-tuning (PEFT) modules of DynaMMo, DAM [21] introduces dynamic aggregated PEFT modules during inference to perform CL. AMM [19] proposes merging convolutional layers to facilitate incremental new class discovery and prevent forgetting fundamental knowledge. Disperse-Then-Merge [48] suggests merging submodels trained on different data partitions during the supervised fine-tuning of LLMs to reduce data bias and mitigate the forgetting of generic pre-trained knowledge.
@@
4.2 Model Merging in Multi-Task/Multi-Objective/Multi-Domain/Auxiliary Learning
@@
In machine learning, to optimize resource efficiency, we typically use a single model to handle multiple tasks, objectives, or data domains with varying distributions. The traditional multi-task learning (MTL), multi-objective learning (MOO), or multi-domain learning (MDL) paradigm requires gathering data from all tasks, objectives, or domains to collaboratively train a model, leading to high data management and model training costs. This approach becomes particularly costly when new tasks, goals, or domains are introduced, as retraining a comprehensive model from scratch using all available data is resource-intensive. Numerous recent studies have proposed efficient methods for integrating knowledge across tasks, goals, or domains by merging models directly.
4.2.1 Knowledge Transfer in Multi-Task Learning
@@
The goal of multi-task learning (MTL) is to enable a single model to perform multiple tasks simultaneously, thereby facilitating knowledge transfer between these tasks [123, 20, 162, 172, 228, 218]. As shown in Figure 1(c), to avoid the high cost of joint training, a straightforward approach is to merge multiple independently trained models on different tasks to accomplish MTL. Almost all of the model merging methods discussed in §2.3 can be used to merge multiple models trained on different tasks to perform MTL. In this section, we take some representative tasks as examples. For MTL tasks in computer vision, Task Arithmetic [71], Ties-Merging [216], AdaMerging [220] and other studies [223, 176, 219] proposed to combine ViT models trained on different visual classification tasks, and the obtained model can complete the object classification of multiple tasks. The results of Task Arithmetic [71] demonstrate that merging independently trained models from any two datasets yields a merged model whose performance is comparable to that of a single-task model. Similarly, ZipIt [170], which merges ResNet architectures trained on different tasks, achieves comparable results. For MTL tasks in natural language processing, DARE [226] introduces a method to assimilate homologous models, augmenting LLMs as a ”free lunch”. For instance, merging WizardLM with WizardMath significantly boosts WizardLM’s performance on GSM8K (a benchmark for evaluating the mathematical reasoning ability of LLMs) from 2.2 to 66.3. Akiba et al. [6] suggest that directly merging an LLM with mathematical capabilities and an LLM with Japanese language proficiency results in a model capable of solving Japanese mathematical problems. Furthermore, numerous studies have demonstrated that combining PEFT modules (such as Adapter, LoRA or Soft-prompts) trained on different tasks can also achieve MTL [180, 234, 11].
4.2.2 Knowledge Transfer in Multi-Objective Optimization
@@
Multi-objective optimization (MOO) aims to optimize multiple objective functions simultaneously. These objective functions may conflict with one another, so the MOO problem typically does not have a single optimal solution. Instead, it involves finding a trade-off among the multiple objectives, which corresponds to identifying a set of Pareto optimal solutions. Tang et al. [178] propose approximating the entire Pareto set using a mixture of experts (MoE) based model merging approach. Specifically, their method trains an independent model for each objective and learns a routing network to balance the trade-offs between the multiple objectives (models). The input of the routing network is the task preference vector, and its output consists of the merging coefficients for the independent models. Considering that directly evaluating Pareto solutions based on the original evaluation metric is time-consuming, MAP [103] proposes a second-order Taylor expansion model as a surrogate model for the true evaluation metric, and further uses an evolutionary algorithm to calculate the Pareto front based on the surrogate model.
4.2.3 Knowledge Transfer in Multi-Domain Learning
@@
Unlike existing model-merging-based MTL approaches that focus on datasets with different object categories, Ye et al. [223] explore model merging across multiple domains, where datasets share the same categories but differ in environmental contexts. To mitigate conflicts between multi-domain models, this paper introduces a weight similarity criterion to assess the correlation between different model layers. For layers with high correlation, a simple weight averaging or RegMean [86] strategy is employed to merge models that have been fine-tuned in different domains of the same task. For layers with low correlation, the weights are flexibly combined using a gating mechanism during the inference phase. Branch-Train-Merge [104] demonstrates the effectiveness of training expert language models on 64 different domains and subsequently merging them.
4.2.4 Knowledge Transfer in Auxiliary Task Learning
@@
The goal of auxiliary task learning (ATL) is to enhance the performance of the target task by leveraging knowledge obtained from related auxiliary tasks. Unlike MTL, which aims to optimize the average performance across all tasks, ATL focuses solely on improving the performance of the main task. However, ATL often encounters the issue of gradient conflict, leading to negative transfer, where the inclusion of auxiliary tasks interferes with the main task’s performance. To mitigate negative transfer, Jiang et al. [84] propose ForkMerge, a method that periodically performs ‘fork’ and ‘merge’ operations. The model is first periodically duplicated into multiple branches: the first branch is trained exclusively on the main task, while the remaining branches are trained jointly on both the main and auxiliary tasks. An optimal merging coefficient is then determined using the validation set to merge the models updated by the various branches. Empirical results show that ForkMerge achieves positive transfer gains across several auxiliary task learning benchmarks.
4.3 Model Merging in Out-of-Distribution/Domain Generalization
@@
4.3.1 Model Merging for Better Out-of-Distribution Generalization
@@
In real-world scenarios, a trained model may be deployed in environments with changing distributions. For example, autonomous driving models are trained on a clean dataset, but in practice, they are vulnerable to unforeseen distributions such as natural corruptions (e.g., camera noise, motion blur) and more significant distribution shifts (e.g., summer to winter) [61, 16]. The goal of OODG is to enhance the model’s ability to generalize to unknown data that significantly differs from the training distribution.
@@
Stochastic weight averaging (SWA) [73] is a straightforward and widely used technique to improve machine learning models’ training stability and OOD performance. From a statistical perspective, weight averaging helps reduce variance during model training. Many works merge intermediate weight states (i.e., checkpoints) from training trajectories while training models [182, 73, 221, 54, 233, 192]. For example, WiSE fine-tuning [209] demonstrates that linearly combining the weights of a pre-trained model and a fine-tuned model can significantly improve accuracy in the case of distribution shifts, while maintaining high accuracy on the original distribution. SWA [73, 54] simply averages all checkpoints from the beginning of a particular epoch to the end of training. This approach is explained to help the model converge to flat rather than sharp local optima, thereby improving generalization [73, 90]. Adaptive SWA [35] highlights that executing SWA too early may lead to underfitting, while executing it too late may result in overfitting. It proposes averaging only when generalization on the validation set improves, effectively combining SWA with an early stopping mechanism. However, simple average weights are often suboptimal. In particular, TWA [106] addresses this by showing that the averaging coefficients of the weights can be determined in a training manner. Consequently, TWA, unlike simple SWA, can perform averaging from the initial epoch of training, eliminating the need to define an additional hyperparameter for the epoch at which weight averaging should start.
@@
In contrast to previous works that average weights obtained along one training trajectory, methods such as Model Soups [208, 245], AdapterSoup [26], Model-Ratatouille [150], WARM [153], WARP [152], PAPA [87], WASH [46], DART [76], and DiWA [149] propose merging multiple independently fine-tuned or trained models. These models are usually more diverse, which improves OOD performance. Independently trained models differ in hyperparameters (e.g., learning rate, weight decay, dropout), batch order, data augmentation techniques (e.g., random crops, horizontal flips), and the number of training steps, among other factors. Specifically, Model-Ratatouille [150], starts from the same initial model, fine-tunes multiple models on an auxiliary task, then continues to fine-tune these models on the target task, and finally merges the diverse models to improve OOD performance. WARM [153] further increases the diversity of fine-tuned models by sampling different checkpoints from the trajectories of the pre-trained model as the initial weights for the downstream preference fine-tuning task. To reduce the additional cost of training multiple models, Model Stock [77] proposes that we can exploit the geometric properties of the weight space and the anchoring effect of pretrained models to approximate the merged weights using only a few fine-tuned models. MEHL-Soup [107] develops a scalable and efficient method to learn model merging coefficients for model soup. It only loads a subset of models for each iteration, significantly reducing the computation and memory requirements of naive model soup for learning merging coefficients.
@@
The above analysis reveals that the SWA lacks diversity due to its reliance on a single trajectory. In contrast, Model Soups and DiWA train independently, which can lead to multiple models with significant differences, resulting in weight averaging failure. To balance these two approaches, Lookaround [233] introduces a gradient descent optimizer based on the weight averaging. This optimizer iteratively performs ‘around’ and ‘average’ steps throughout the optimization process. In the ‘around’ step, multiple independent models are trained from the same starting point, each using different data augmentations. In the ‘average’ step, the diverse models are averaged, and the result is used as the starting point for the next iteration.
4.3.2 Model Merging for Better Domain Generalization
Domain generalization methods aim to generalize to an unknown target domain using only training data from source domains. For instance, in the context of traffic sign recognition, the training data for a machine learning (ML) model tasked with identifying traffic signs in various urban environments come from multiple cities (i.e., source domains). However, when deployed, the model must recognize traffic signs in new urban environments (i.e., target domains) that it has never encountered before. Existing DG methods can be classified into domain alignment, data augmentation, regularization, and meta-learning frameworks [8]. Complementary to these approaches, model merging techniques can be seamlessly integrated to further improve out-of-domain performance without modification [17, 8, 109]. Specifically, model merging in DG mainly occurs during the training process of the source domain model. Merging the intermediate weight states from different training stages helps improve the stability and generalization of the final model.

### v5

@@ 5. Application of Model Merging in Different Machine Learning Subfields
Table 5. A summary of the application of model merging techniques in different machine learning subfields.
@@
5.1. Model Merging in Continual Learning
@@
Continual Learning (CL) involves training a model using a streaming, non-stationary data stream. The primary challenge in CL is the ‘catastrophic forgetting’ (Robins, 1995; Wang et al., 2023b) problem; that is, the CL model’s prediction accuracy for old tasks drops dramatically after training on new tasks. The mainstream CL methods are mainly divided into memory replay-based methods, architecture expansion-based methods, regularization-based methods, and subspace projection-based methods (Wang et al., 2023b). Recently, there has been a growing interest in using model merging to address the catastrophic forgetting problem. This novel approach offers several benefits, such as avoiding additional parameters and inference costs associated with network expansion-based methods and eliminating the need to cache old data as required by memory-based methods.
@@
To overcome catastrophic forgetting and maintain new learning ability, many works have proposed to merge models for old tasks and update models for new tasks (Wortsman et al., 2022b; Schumann et al., 2024; Panda et al., 2024). For example, WiSE-FT (Wortsman et al., 2022b) shows that merging the fine-tuned model with the pretrained model can effectively mitigate the forgetting of general knowledge. Tangent Model Composition (Liu and Soatto, 2023) proposes fine-tuning each task independently in the tangent space of the pre-trained model and then merging the resulting models to perform CL. In addition, ITA (Porrello et al., 2024) emphasizes the necessity for the fine-tuned model to be in the same basin as the pre-trained model to ensure the composability of nonlinear models. It introduces a regularization term similar to EWC (Kirkpatrick et al., 2017) in traditional CL to constrain the distance between the fine-tuned weights and the pre-trained weights when training the independent model. WARP (Ramé et al., 2024a) suggests linearly interpolating the pre-trained LLM’s weights with its aligned weights via RLHF on a preference dataset, thus mitigating the forgetting of knowledge from the pre-trained LLM. BAM (Alexandrov et al., 2024) continuously adapts LLMs to new languages by merging models while preserving general capabilities. Model Tailor (Zhu et al., 2024) explores the problem of catastrophic forgetting during fine-tuning of MLLMs, and proposes to merge only the most important subset of parameters in the fine-tuned MLLM model into the pre-trained MLLM model, so as to retain the generalization ability of the pre-trained model as much as possible, while compensating the selected weights to reduce the performance of the fine-tuning task. MagMax (Marczak et al., 2024) merges pruned task vectors to further alleviate parameter sign conflicts and old knowledge forgetting. Equifinality, PAINT (Ilharco et al., 2022) and LM-Cocktail (Xiao et al., 2024) interpolate the weights of the fine-tuned model and the zero-shot model to improve accuracy on downstream tasks without degrading accuracy on supported/general tasks.
@@
In contrast to merging full models, some research focuses on merging parameter-efficient modules. Chitale et al. (2023) propose a CL method based on task arithmetic (Ilharco et al., 2023). This method first fine-tunes a task-specific LoRA for each task, then constructs a task vector based on the difference between fine-tuned and pre-trained models. Multiple task vectors are then merged, and a small amount of data (10 samples per class) is used to fine-tune the merged model. Compared to traditional CL methods, particularly those based on replay, this approach eliminates the need to replay data from old tasks at each iteration, thereby accelerating model training. Additionally, fine-tuning the merged model with a class-balanced subset helps mitigate CL model bias. Similarly, DynaMMo (Qazi et al., 2024) applies lightweight model merging (i.e., Adapter) in a CL setting for medical images. In contrast to architecture expansion-based CL methods, this approach does not result in a linear increase in the number of parameters with the number of tasks. Unlike the static aggregated parameter-efficient fine-tuning (PEFT) modules of DynaMMo, DAM (Cheng et al., 2024) introduces dynamic aggregated PEFT modules during inference to perform CL. AMM (Chen et al., 2024b) proposes merging convolutional layers to facilitate incremental new class discovery and prevent forgetting fundamental knowledge. Disperse-Then-Merge (Fu et al., 2024) suggests merging submodels trained on different data partitions during the supervised fine-tuning of LLMs to reduce data bias and mitigate the forgetting of generic pre-trained knowledge.
@@
5.2. Model Merging in Multi-Task/Multi-Objective/Multi-Domain/Auxiliary Learning
@@
In machine learning, to optimize resource efficiency, we typically use a single model to handle multiple tasks, objectives, or domains with varying distributions. The traditional multi-task learning (MTL), multi-objective learning (MOO), or multi-domain learning (MDL) paradigm requires gathering data from all tasks, objectives, or domains to collaboratively train a model, leading to high data management and model training costs. This approach becomes particularly costly when new tasks, goals, or domains are introduced, as retraining a comprehensive model from scratch using all available data is resource-intensive. Numerous recent studies have proposed efficient methods for integrating knowledge across tasks, goals, or domains by merging models directly.
5.2.1. Knowledge Transfer in Multi-Task Learning
@@
The goal of MTL is to enable a single model to perform multiple tasks simultaneously, thereby facilitating knowledge transfer between these tasks (Ma et al., 2018; Chen et al., 2018; Sener and Koltun, 2018; Sun et al., 2020; Yu et al., 2020; Yang et al., 2023a). As shown in Figure 1(b), to avoid the high cost of joint training, a straightforward approach is to merge multiple independently trained models on different tasks to accomplish MTL. Almost all of the model merging methods discussed in §2.3 can be used to merge multiple models trained on different tasks to perform MTL. In this section, we take some representative tasks as examples. For MTL tasks in computer vision, Task Arithmetic (Ilharco et al., 2023), Ties-Merging (Yadav et al., 2023), AdaMerging (Yang et al., 2024c) and other studies (Ye et al., 2023; Tang et al., 2023; Yang et al., 2024a) proposed to combine ViT models trained on different visual classification tasks, and the obtained model can complete the object classification of multiple tasks. The results of Task Arithmetic (Ilharco et al., 2023) demonstrate that merging independently trained models from any two datasets yields a merged model whose performance is comparable to that of a single-task model. Similarly, ZipIt (Stoica et al., 2024a), which merges ResNet architectures trained on different tasks, achieves comparable results. For MTL tasks in natural language processing, DARE (Yu et al., 2024b) introduces a method to assimilate homologous models, augmenting LLMs as a ”free lunch”. For instance, merging WizardLM with WizardMath significantly boosts WizardLM’s performance on GSM8K (a benchmark for evaluating the mathematical reasoning ability of LLMs) from 2.2 to 66.3. Akiba et al. (2025) suggest that directly merging an LLM with mathematical capabilities and an LLM with Japanese language proficiency results in a model capable of solving Japanese mathematical problems. Furthermore, numerous studies have demonstrated that combining PEFT modules trained on different tasks can also achieve MTL (Tang et al., 2024c; Zhang et al., 2023a; Belanec et al., 2024).
5.2.2. Knowledge Transfer in Multi-Objective Optimization
@@
MOO aims to optimize multiple objective functions simultaneously. These objective functions may conflict with one another, so the MOO problem typically does not have a single optimal solution. Instead, it involves finding a trade-off among the multiple objectives, which corresponds to identifying a set of Pareto optimal solutions. Tang et al. (2024a) propose approximating the entire Pareto set using a mixture of experts based model merging approach. Specifically, their method trains an independent model for each objective and learns a routing network to balance the trade-offs between the multiple objectives. The input of the routing network is the task preference vector, and its output consists of the merging coefficients for the independent models. Pareto Merging (Chen and Kwok, 2025) also formulates different preferences as an MOO problem, and generates a Pareto set of merged models. Considering that directly evaluating Pareto solutions based on the original evaluation metric is time-consuming, MAP (Li et al., 2024d) proposes a second-order Taylor expansion model as a surrogate model for the true evaluation metric, and further uses an evolutionary algorithm to calculate the Pareto front based on the surrogate model.
5.2.3. Knowledge Transfer in Multi-Domain Learning
@@
Unlike existing model-merging-based MTL approaches that focus on datasets with different object categories, Ye et al. (2023) explore model merging across multiple domains, where datasets share the same categories but differ in environmental contexts. To mitigate conflicts between multi-domain models, this paper introduces a weight similarity criterion to assess the correlation between different model layers. For layers with high correlation, a simple weight averaging or RegMean (Jin et al., 2023) strategy is employed to merge models that have been fine-tuned in different domains of the same task. For layers with low correlation, the weights are flexibly combined using a gating mechanism during the inference phase. Branch-Train-Merge (Li et al., 2022) demonstrates the effectiveness of training expert language models on 64 different domains and subsequently merging them.
5.2.4. Knowledge Transfer in Auxiliary Task Learning
@@
The goal of ATL is to enhance the performance of the target task by leveraging knowledge obtained from related auxiliary tasks. Unlike MTL, which aims to optimize the average performance across all tasks, ATL focuses solely on improving the performance of the main task. However, ATL often encounters the issue of gradient conflict, leading to negative transfer, where the inclusion of auxiliary tasks interferes with the main task’s performance. To mitigate negative transfer, Jiang et al. (2023) propose ForkMerge, which periodically performs ‘fork’ and ‘merge’ operations. The model is first periodically duplicated into multiple branches: the first branch is trained exclusively on the main task, while the remaining branches are trained jointly on both the main and auxiliary tasks. An optimal merging coefficient is then determined using the validation set to merge the models updated by the various branches. Empirical results show that ForkMerge achieves positive transfer gains across several ATL benchmarks.
5.3. Model Merging in Out-of-Distribution/Domain Generalization
@@
5.3.1. Model Merging for Better Out-of-Distribution Generalization
@@
In real-world scenarios, a trained model may be deployed in environments with changing distributions. For example, autonomous driving models are trained on a clean dataset, but in practice, they are vulnerable to unforeseen distributions such as natural corruptions (e.g., camera noise, motion blur) and more significant distribution shifts (e.g., summer to winter) (Hendrycks and Dietterich, 2019; Cai et al., 2023). The goal of OODG is to enhance the model’s ability to generalize to unknown data that significantly differs from the training distribution.
@@
Stochastic weight averaging (SWA) (Izmailov et al., 2018) is a straightforward and widely used technique to improve machine learning models’ training stability and OOD performance. From a statistical perspective, weight averaging helps reduce variance during model training. Many works merge intermediate weight states (i.e., checkpoints) from training trajectories while training models (Tarvainen and Valpola, 2017; Izmailov et al., 2018; Yang et al., 2019; Gupta et al., 2020; Zhang et al., 2023b; von Oswald et al., 2021). For example, WiSE fine-tuning (Wortsman et al., 2022b) demonstrates that linearly combining the weights of a pre-trained model and a fine-tuned model can significantly improve accuracy in the case of distribution shifts, while maintaining high accuracy on the original distribution. SWA (Izmailov et al., 2018; Gupta et al., 2020) simply averages all checkpoints from the beginning of a particular epoch to the end of training. This approach is explained to help the model converge to flat rather than sharp local optima, thereby improving generalization (Izmailov et al., 2018; Kaddour et al., 2022). Adaptive SWA (Demir et al., 2024) highlights that executing SWA too early may lead to underfitting, while executing it too late may result in overfitting. It proposes averaging only when generalization on the validation set improves, effectively combining SWA with an early stopping mechanism. TWA (Li et al., 2023a) addresses this by showing that the averaging coefficients of the weights can be determined in a training manner. Consequently, TWA, unlike simple SWA, can perform averaging from the initial epoch of training, eliminating the need to define an additional hyperparameter for the epoch at which weight averaging should start.
@@
In contrast to previous works that average weights obtained along one training trajectory, methods such as Model Soups (Wortsman et al., 2022a; Zimmer et al., 2024), AdapterSoup (Chronopoulou et al., 2023a), Model-Ratatouille (Ramé et al., 2023), WARM (Ramé et al., 2024b), WARP (Ramé et al., 2024a), PAPA (Jolicoeur-Martineau et al., 2024), WASH (Fournier et al., 2024), DART (Jain et al., 2023), and DiWA (Rame et al., 2022) propose merging multiple independently fine-tuned or trained models. These models are usually more diverse, which improves OOD performance. Independently trained models differ in hyperparameters (e.g., learning rate, weight decay, dropout), batch order, data augmentation techniques (e.g., random crops, horizontal flips), and the number of training steps, among other factors. Specifically, Model-Ratatouille (Ramé et al., 2023), starts from the same initial model, fine-tunes multiple models on an auxiliary task, then continues to fine-tune these models on the target task, and finally merges the diverse models to improve OOD performance. WARM (Ramé et al., 2024b) further increases the diversity of fine-tuned models by sampling different checkpoints from the trajectories of the pre-trained model as the initial weights for the downstream preference fine-tuning task. To reduce the additional cost of training multiple models, Model Stock (Jang et al., 2024) proposes that we can exploit the geometric properties of the weight space and the anchoring effect of pretrained models to approximate the merged weights using only a few fine-tuned models. MEHL-Soup (Li et al., 2024c) develops a scalable and efficient method to learn merging coefficients for model soup. It only loads a subset of models for each iteration, significantly reducing the computation and memory requirements of naive model soup for learning merging coefficients.
@@
The above analysis reveals that the SWA lacks diversity due to its reliance on a single trajectory. In contrast, Model Soups and DiWA train independently, which can lead to multiple models with significant differences, resulting in weight averaging failure. To balance these two approaches, Lookaround (Zhang et al., 2023b) introduces a gradient descent optimizer based on the concept of weight averaging. This optimizer iteratively performs ‘around’ and ‘average’ steps throughout the optimization process. In the ‘around’ step, multiple independent models are trained from the same starting point, each using different data augmentations. In the ‘average’ step, the diverse models are averaged, and the result is used as the starting point for the next iteration.
5.3.2. Model Merging for Better Domain Generalization